the sum of three consecutive odd intergers is 165. what are the integers?
the ages od edna, ellie, & elsa are consecutive intergers the sum of their ages is 120.

Answer by oberobic(2304)

1) Sum of 3 consecutive odd integers is 165.
.
You could use x, y, and z; but then you need 3 equations to solve the 3 unknowns.
But we are told they a consecutive odd integers, so we can use x, x+2, and x+4.
.
x + x+2 + x+4 = 165
3x + 6 = 165
3x = 159
x = 53
so
x+2 =55
x+4 = 57
.
Answer: 53, 55, and 57
.
2) Sum of consecutive integers is 120
.
x + x+1 + x+2 = 120
3x +3 = 120
3x = 117
x = 39
.
x+1 = 40
x+2 = 41
.
Answer: 39, 40, and 41
.
Done.